# Disentangling Exploration of Large Language Models by Optimal Exploitation

### v1

@@ Abstract
Exploration is a crucial skill for self-improvement and open-ended problem-solving. However, it remains uncertain whether large language models can effectively explore the state-space. Existing evaluations predominantly focus on the trade-off between exploration and exploitation, often assessed in multi-armed bandit problems. In contrast, this work isolates exploration as the sole objective, tasking the agent with delivering information that enhances future returns. For the evaluation, we propose to decompose missing rewards into exploration and exploitation components by measuring the optimal achievable return for the states already explored. Our experiments with various LLMs reveal that most models struggle to sufficiently explore the state-space and that weak exploration is insufficient. We observe a positive correlation between model size and exploration performance, with larger models demonstrating superior capabilities. Furthermore, we show that our decomposition provides insights into differences in behaviors driven by agent instructions during prompt engineering, offering a valuable tool for refining LLM performance in exploratory tasks.
@@ 1 Introduction
Large Language Models as decision making agents with a good inductive bias.    Recently, large language models (LLMs) have demonstrated promising results in various decision making tasks such as web browsing (Yao et al., 2022; Shinn et al., 2024; Ma et al., 2023), game-playing (Paglieri et al., 2024), and tasks in simulated households (Yao et al., 2022; Shinn et al., 2024). This way, LLMs act as agents that observe states and take actions in different environments. Through their vast internal knowledge-base and autoregressive in-context reasoning capabilities, the models are supposed to quickly adapt to new tasks. However, previous work has shown that LLMs struggle with solving increasingly complex environments due to several limitations: For example, the ability to learn from mistakes is often limited (Huang et al., 2023) and LLMs have difficulties with planning over long horizons (Kambhampati et al., 2024). The examples emphasize that understanding LLM abilities is essential for their risk assessment in real life applications, and future development.
@@
Intelligent exploration is essential for self-improvement.    For our human nature, intelligent exploration is an essential skill for improvement in many tasks. Exploration is a critical process for systematically seeking new information or testing novel actions to reduce uncertainty and improve decision-making over time. Traditional approaches often rely on stochastic noise (Mnih, 2013; Lillicrap, 2015) or reward-shaping (Bellemare et al., 2016). Such methods induce redundancy and differ from human decision making which is mainly driven by reasoning about current information, past steps, and the unknown potential of yet unknown actions. Therefore, recent works proposed several approaches towards more intelligent exploration (Huang et al., 2024; Lu et al., 2024; Nie et al., 2024). Hereby, due to their emergent abilities (Wei et al., 2022), LLMs could have the potential to revolutionize discovery by identifying patterns, leveraging prior knowledge, and intelligently navigating the state space. LLMs may transform exploration into a targeted and efficient process.
Limited focus on exploration as an independent capability. While exploration is acknowledged as a crucial component for learning and decision making, prior work has predominantly investigated it in conjunction with exploitation (Krishnamurthy et al., 2024; Nie et al., 2024; Paglieri et al., 2024; Ke et al., 2024). These studies primarily investigate the exploration-exploitation trade-off by comparing LLM-based agents to upper confidence bound algorithms, rather than examining the potential of agents dedicated solely to systematic state-space traversal. In contrast to these approaches, Huang et al. (2024) propose decoupling the exploration process as a separate agent. Despite this advancement, the work lacks an evaluation centered on exploration, leaving the full-potential of the decoupled agent unexplored. At the same time, new environments that challenge LLMs with tasks requiring meaningful state-space exploration are needed. Domains like ALFWorld (Shridhar et al., 2020) often feature straightforward cause-and-effect relationships, which do not demand robust or long-horizon exploration. In contrast, traditional reinforcement learning benchmarks like the NetHack Learning Environment (Küttler et al., 2020) complicate the assessment by emphasizing navigation deficiencies in LLMs (Paglieri et al., 2024), rather than providing opportunities to evaluate intelligent and abstract strategies.
Contribution.    In this paper, we claim that exploration will be an essential skill for future agents. Recent studies (Krishnamurthy et al., 2024; Nie et al., 2024) found LLMs balancing exploration and exploitation often struggle to discover the best actions. Therefore, we deliberately prompt the LLM to focus on exploration. Hence, isolating the this behavior. In the controlled setting, we argue that conventional metrics, such as agent returns, LLM exploitation, or the convergence towards the best sequence of actions, (Huang et al., 2024; Krishnamurthy et al., 2024) are insufficient. To fill this gap, we measure the optimal achievable exploitation return. Contrary to prior work (Krishnamurthy et al., 2024) on multi-armed bandits, our evaluation framework is designed to be environment-agnostic, making it applicable to a broader range of tasks and scenarios in future research. Using the optimal exploitation return, we decompose missing rewards into exploration and exploitation components. This decomposition measures exploration progress more accurately and enables insights into reward sacrification. To demonstrate the utility of our approach, we conduct a comprehensive evaluation of various popular open-source and closed LLMs, leveraging an adapted version of Dark Treasure Room (Norman & Clune, 2023). The adaptation is tailored for high-level and semantic action spaces and introduces a complex exploration challenge that cannot be resolved through reliance on internal world knowledge alone. To the best of our knowledge, this work is the first to evaluate and compare the state-space exploration capabilities of different LLMs.
@@
Results.    Our findings indicate that most LLMs face challenges with state-space exploration, when the task is isolated. These results align with prior work (Krishnamurthy et al., 2024; Paglieri et al., 2024), which highlights the trade-off between exploration and exploitation. We observe that state-space coverage diminishes significantly when agents are required to plan over extended horizons, indicating a potential limitation in long-term exploratory capabilities. Contrary to earlier claims (Huang et al., 2024), our analysis demonstrates that weak LLMs are insufficient. Notably, we identified a significant trend wherein stronger exploration capabilities correlate with increased model scale, suggesting that future, more advanced LLMs may possess enhanced exploration potential. Our method further uncovered model-specific differences in reward sacrification and we demonstrated how the impact of agent instructions on exploratory behavior can be assessed, providing insights into the design of prompts during prompt-engineering.
@@ 2 Related Work
Recently, exploring the capabilities of LLMs gained significant attention and has been investigated from different perspectives. Prior works include, but are not limited to, planning (Kambhampati et al., 2024; Song et al., 2023; Valmeekam et al., 2023), world models (Hao et al., 2023; Guan et al., 2023), logical- (Xu et al., 2023; Parmar et al., 2024), commonsense- (Li et al., 2021), and mathematical reasoning (Imani et al., 2023; Yuan et al., 2023). Our research focuses on intelligent exploration of the state-space in sequential decision making. Generally, we summarize existing methods into three streams of work:
@@
Large Language Models in decision making.    A broad line of works has applied LLMs in decision making. Relatively few specifically addressed the challenge of exploration. The majority of existing research focuses on multi-armed bandit problems (Park et al., 2024; Krishnamurthy et al., 2024; Nie et al., 2024). For example, (Krishnamurthy et al., 2024) investigate scenarios involving multiple buttons with underlying stochastic rewards. The LLM is presented past actions with their rewards and must maximize its return over multiple trials. However, multi-armed bandits lack the rich sequential state-space of many real-world applications. Other research examined exploration in hypothesis testing (Piriyakulkij et al., 2024; Ke et al., 2024). Nevertheless, only a limited number of works (Ma et al., 2023; Paul, 2023; Huang et al., 2024; Lu et al., 2024) proposed methods to enhance state-space traversal. Huang et al. (2024) argue that smaller-scale models can adequately handle exploration, reserving larger, more competent models for exploitation. While their approach involves prompting LLMs to explore as as a system component, it does not provide an isolated evaluation. Contrary, the focus of our study is on an evaluation of the exploration capabilities.
@@
Exploration in reinforcement learning.    Our evaluation is closely related to exploration in reinforcement learning (RL). Early approaches predominantly relied on random methods, such as epsilon-greedy strategies (Mnih, 2013; Lillicrap, 2015). More recent research has shifted toward developing intelligent exploration techniques (Zhang et al., 2021; Ecoffet et al., 2019; Norman & Clune, 2023; Mazoure et al., 2023). Methods that decouple exploration and exploitation are particularly relevant to our work (Liu et al., 2021; Avner et al., 2012; Norman & Clune, 2023; Zhang et al., 2021; Duan et al., 2016). In methods, such as First-Explore (Norman & Clune, 2023) exploration and exploitation are handled by separate policies, with the exploration policy informing subsequent exploitation decisions. Generally, RL methods focus on stabilizing learning of the decoupled policies. Evaluations use non-optimal exploitation returns as optimal policies can be hard to obtain in complex and high-dimensional spaces. Contrary, we measure the optimal exploitation return for the evaluation of LLMs, as LLM exploitation and agent returns do not provide reliable estimates.
@@
Evaluation Environments.    A variety of environments have been developed to assess different capabilities of LLMs. Notable examples that require exploration to solve tasks include TextWorld (Côté et al., 2019), Minecraft (Fan et al., 2022), Crafter (Hafner, 2021), NetHack (Küttler et al., 2020), MiniHack (Samvelyan et al., 2021), MiniGrid (Chevalier-Boisvert et al., 2024), and BabyAI (Chevalier-Boisvert et al., 2018). Additionally, there is growing interest in environments designed specifically for web agents, such as WebArena (Zhou et al., 2023). Comprehensive benchmarks, including SmartPlay (Wu et al., 2023), AgentBench (Liu et al., 2023), and BALROG (Paglieri et al., 2024), aggregate multiple environments to evaluate agents across a wide range of capabilities. BALROG has documented issues such as agents frequently exhibiting aimless movement and redundant behaviors, highlighting challenges in effective exploration. We investigate exploration using an adapted version of the Dark Treasure Room environment (Norman & Clune, 2023). The difference in our work lies in the isolation of the exploration. Our environment also often differs in limiting prior knowledge, providing relatively dense rewards, and maintaining a high-level action space.
3 Evaluating Large Language Models for Exploration
In the following section, we introduce our evaluation approach and environment tailored for LLM exploration capabilities in sequential decision-making. Importantly, our method is environment-agnostic and independent of specific prompts. Future research can extend this framework to facilitate comparisons of agent performance across diverse environments and contexts.
@@ 3.1 Theoretical Framework
Deterministic Markov Decision Process    We study a deterministic Markov decision process $(\mathcal{S},\mathcal{A},\mathcal{P},r,\gamma)$ where $\mathcal{S}$ is a finite state space, $\mathcal{A}$ is the finite action space, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ defines the transition dynamics of the environment, $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function and $\gamma$ the discount factor. At each time step $t$, an agent samples an action from policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$ based on the current observation $s_{t}\in\mathcal{S}$ and executes it in the environment. The environment transitions and the agent receives a reward $r_{t}$. In an episodic MDP, the procedure repeats until a terminal state $s_{T}$ is reached or the maximum number of steps exceeds. The sequence of state, action, and reward triples until $s_{T}$ is called a trajectory and describes an episode $\tau_{i}=\left(s_{0},a_{0},r_{0},s_{1},a_{1},r_{1},s_{2},\dots\right)$. In our setting, the agent has access to the current and history of prior trajectories $h_{i}=(\tau_{1},\dots,\tau_{i})$ in episode $i$. The goal is to find the sequence of actions maximizing the cumulative return $R_{i}=\sum_{t=1}^{T}r_{t}$.
@@
Exploitation versus Exploration.    The cumulative return $R^{\text{LLM}}_{i}$ of an agent is influenced by its ability to balance exploitation and exploration. Exploitation involves leveraging past high-rewarding actions to maximize immediate returns based on the agent’s current knowledge. In contrast, exploration emphasizes interacting with the environment to gather new information, which can enhance future decision-making and returns. When the agent is fully dedicated to exploitation, it leverages the information in past trajectories $h_{i}$ to estimate the optimal action sequence $\pi_{\text{exploit}}^{*}(h_{i})=\arg\max_{\pi}\sum_{t=0}^{T}r(s_{t},\pi(s_{t}))$. Conversely, an agent focused on exploration $\pi^{*}_{\text{explore}}$ seeks to augment the trajectory history $h_{i}$ by incorporating new trajectories that maximize the potential future exploitation return. Under the assumption of an optimal exploitation return $R^{\text{exploit}}_{i}$, which ignores redundant or negatively impactful trajectories and whose return monotonically increases over episodes, the objective of the exploration policy is to maximize the exploitation return at the end of an exploration $\pi_{\text{explore}}^{*}=\arg\max_{\pi_{\text{explore}}}\mathbb{E}\left[R^{\text{exploit}}_{\text{final}}\mid h_{i}=h_{i-1}\cup\tau_{\text{explore}}\right]$.
@@
The need for optimal exploitation.    When agents do both, exploitation and exploration, progress can be evaluated by temporarily disabling exploration and measuring the agent’s return. However, for an agent fully dedicated to exploration, its return alone does not accurately reflect success (see figures 2(a), 2(b)), as successful exploration does not directly result in rewards. Also measuring exploitation returns with an LLM is similarly hard to exploration and does not give reliable estimates (appendix B.1). However, due to the simplistic nature of LLM environments, we can measure the optimal exploitation return $R^{\text{exploit}}_{i}$ by using, for example, dynamic programming methods which guarantee the optimal exploitation based on history $h_{i}$. By calculating the optimal exploitation return, we can assess an algorithm’s exploration capabilities more accurately without conflating them with the exploitation algorithm.
@@
Decomposing the total return gap.    Based on the optimal exploitation return, we can decompose the total return gap, i.e., the difference of the maximal achievable return of the environment and the actual achieved return. Let $R^{max}$ denote the maximal achievable reward of the environment; it is independent of the agent and can be predefined for evaluation purposes. As previously, let $R^{LLM}_{i}$ be the agent reward and $R^{exploit}_{i}$ the optimal exploitation reward by acting according to policy $\pi^{*}_{\text{exploit}}$. Contrary to prior evaluations, $\pi^{*}_{\text{exploit}}$ is not determined by an agent. Note that it follows $R^{max}\leq R^{exploit}\leq R^{LLM}$.
We let $\Delta^{total}$ denote the total reward gap, which is calculated as the difference of maximal achievable and the actual achieved return by the agent:
@@
The total return gap decomposes in the the exploration and exploitation gap.
3.2 Environment
(a) 4x4
(b) 5x5
(c) 7x7
(d) LLM Agent
Figure 1: Our variant of Dark Treasure Room. The agent navigates a grid of rooms by interacting with doors and rewarding balls. The agent must select the actions that collect most rewards.
Decision making in room worlds.    We employ grids of interconnected rooms, inspired by traditional grid worlds commonly used in RL. We use a higher-level action space that is conceptually similar to TextWorld (Côté et al., 2019) environments and described in appendix A.3. However, our domain features only minimal semantic associations. Each grid has a fixed layout and starting point. Episodes terminate either when the agent collects three objects or exceeds a maximum number of door traversals. This number is calibrated to the distance between the agent’s starting position and the furthest room in the environment, ensuring that the exploration budget appropriately scales with the complexity of the environment. At the beginning of each episode, the agent is reset to its starting position, requiring it to strategically plan exploration over extended horizons.
Dark Treasure Rooms.    The task is an adapted version of Dark Treasure Room (Norman & Clune, 2023) that is more aligned with the capabilities of LLMs. With our room worlds, we replace the low-level spatial navigation with high-level object interactions. Each map contains balls placed randomly, with rewards for collecting each ball sampled from a uniform distribution $U(1,10)$. The agent’s goal is to maximize the return by collecting the three highest-reward balls. As ball locations and their associated rewards are randomly generated, the optimal exploration strategy, across multiple layouts, requires full coverage of the state space to locate all potential rewards and reduce uncertainty. Inadequate exploration risks the exploitation collecting suboptimal rewards due to incomplete information about ball locations. Nevertheless, our evaluation method can also work with environments not requiring full state-space coverage in future work. We evaluate the agents across three grid sizes: $4\times 4$, $5\times 5$, and $\times 7$. These configurations are designed to test the scalability of the agent’s planning horizon. Exemplary environments and an agent interaction are depicted in Figure 1. See appendix A for a full list of our grids.

### v2

@@ Abstract
Exploration is a crucial skill for self-improvement and open-ended problem-solving. However, it remains unclear if large language models can effectively explore the state-space within an unknown environment. This work isolates exploration as the sole objective, tasking the agent with delivering information that enhances future returns. Within this framework, we argue that measuring agent returns is not sufficient for a fair evaluation and decompose missing rewards into exploration and exploitation components based on the optimal achievable return. Comprehensive experiments with various models reveal that most struggle to sufficiently explore the state-space and weak exploration is insufficient. We observe a positive correlation between parameter count and exploration performance, with larger models demonstrating superior capabilities. Furthermore, we show that our decomposition provides insights into differences in behaviors driven by prompt engineering, offering a valuable tool for refining performance in exploratory tasks.
@@ 1 Introduction
Large language models in decision making. Recently, large language models (LLMs) have demonstrated promising results in various decision making tasks such as web browsing Yao et al. (2022b); Shinn et al. (2024); Ma et al. (2023), game-playing Paglieri et al. (2024), and simulated households Yao et al. (2022b); Shinn et al. (2024). Hereby, LLMs act as agents that observe states and take actions in environments. Through their vast internal knowledge-base and autoregressive in-context reasoning capabilities, the models are supposed to quickly adapt to new tasks. However, previous work has shown that LLMs struggle with solving complex problems due to several shortcomings: For example, the ability to learn from mistakes is often limited Huang et al. (2023) and LLMs have difficulties with planning over long horizons Kambhampati et al. (2024).
@@
Exploration is essential for self-improvement. Intelligent exploration is an important skill for improvement in many tasks. Hereby, exploration is a critical process for systematically seeking novel information to reduce uncertainty and improve decision-making over time. Traditional approaches often rely on stochastic noise Mnih (2013); Lillicrap (2015) or reward-shaping Bellemare et al. (2016). Such methods induce redundancy and differ from human decision making which is mainly driven by reasoning about current information, past steps, and the unknown potential of yet unknown actions. Therefore, recent works proposed several approaches towards more intelligent exploration Huang et al. (2024); Lu et al. (2024); Nie et al. (2024). Hereby, LLMs could have the potential to assist discovery by identifying patterns, leveraging prior knowledge, and intelligently navigating the state-space. LLMs may transform exploration into a targeted and efficient process.
Limited focus on exploration as an independent ability. While exploration is acknowledged as a crucial component for learning and LLM decision making, prior work has predominantly investigated it in conjunction with exploitation Krishnamurthy et al. (2024); Nie et al. (2024); Paglieri et al. (2024); Ke et al. (2024); Chen et al. (2024). In this context, exploration is often measured indirectly through cumulative rewards or success rates. However, these approaches conflate the ability with overall performance. It is difficult to isolate an agent’s exploration, as true progress is independent of the agent’s return. While seeking information, agents are supposed to sacrifice short-term rewards in favor of long-term understanding. Also, strong exploration does not always translate into high performance if the agent lacks the ability to exploit the knowledge it has gathered. An agent may thoroughly explore an environment but fail to leverage its findings. Therefore, it is essential to develop methods that accurately measure exploration of LLMs, as it helps to foster our understanding of their learning and adaptability in sequential decision making.
Figure 1: An overview of our evaluation procedure based on an optimal exploitation. The agent navigates a maze or grid of rooms by interacting with doors and balls. Overall, the goal is to collect information about high rewarding balls. After each episode, we calculate the optimal sequence of actions to measure exploration progress based on the agent’s current memory.
Contribution. In this paper, we claim that exploration will be an essential skill for future agents. Recent studies Krishnamurthy et al. (2024); Nie et al. (2024) investigated LLMs balancing exploration and exploitation where they found that they often struggle to achieve high rewards. Conversely, we disentangle exploration for an independent evaluation. In this controlled setting, we argue that conventional metrics, such as agent returns, LLM exploitation, or the convergence towards the best sequence of actions Huang et al. (2024); Krishnamurthy et al. (2024) are insufficient. To fill this gap, we make the following contributions:
  • We propose to measure the optimal exploitation return (see Figure 1) and decompose missing rewards into their exploration and exploitation components. This decomposition measures exploration progress accurately and enables insights into reward sacrification.
  • To the best of our knowledge, we are the first to conduct a comprehensive evaluation of various popular open-source and closed LLMs with the exploration being disentangled.
Our evaluation framework is designed to be environment-agnostic, making it applicable to a broader range of tasks and scenarios in future research.
@@
Results. Our findings indicate that most LLMs face challenges with state-space exploration, when the task is isolated. These results align with prior work Krishnamurthy et al. (2024); Paglieri et al. (2024), which highlight the trade-off between exploration and exploitation. We observe that state-space coverage diminishes significantly when agents are required to plan over extended horizons, indicating a potential limitation in long-term exploratory capabilities. Contrary to earlier claims Huang et al. (2024), our analysis demonstrates that weak LLMs are insufficient. Notably, we identified a significant trend wherein stronger exploration capabilities correlate with increased model scale, suggesting that future, more advanced LLMs may possess enhanced exploration potential. Our method further uncovered model-specific differences in reward sacrification and we demonstrated how the impact of agent instructions on exploratory behavior can be assessed, providing insights into the design of prompts during prompt-engineering.
@@ 2 Related Work
Recently, exploring the capabilities of LLMs gained significant attention and has been investigated from different perspectives. Prior works include, but are not limited to aspects such as planning Kambhampati et al. (2024); Song et al. (2023); Valmeekam et al. (2023), world models Hao et al. (2023); Guan et al. (2023), logical- Xu et al. (2023); Parmar et al. (2024), commonsense- Li et al. (2021), and mathematical reasoning Imani et al. (2023); Yuan et al. (2023). Our research focuses on intelligent exploration of the state-space in sequential decision making. Generally, we summarize existing methods into three streams of work:
@@
Large language models in decision making. A broad line of works has applied LLMs in decision making. Relatively few addressed the challenge of exploration. The majority of existing research focuses on multi-armed bandit problems Park et al. (2024); Krishnamurthy et al. (2024); Nie et al. (2024); Chen et al. (2024). For example, Krishnamurthy et al. (2024) analyze scenarios involving multiple buttons with underlying stochastic rewards. The LLM is presented past actions with their rewards and must maximize the return over multiple trials. However, multi-armed bandits lack the rich sequential state-space of many real-world applications. Other research examined exploration in hypothesis testing Piriyakulkij et al. (2024); Ke et al. (2024). Only a limited number of works Ma et al. (2023); Paul (2023); Huang et al. (2024); Lu et al. (2024) proposed methods for sequential state-spaces. Lu et al. (2024) suggest returning to interesting states. Huang et al. (2024) argue that smaller-scale models can handle exploration, reserving larger models for exploitation. Generally, evaluations mostly focus on agent returns. While exploration LLMs are used as system components, an isolated evaluation is missing. Therefore, the focus of our study is specifically on LLMs that gather information about their environment.
@@
Exploration in reinforcement learning. Our evaluation is closely related to exploration in RL. Early approaches predominantly relied on random methods, such as epsilon-greedy strategies Mnih (2013); Lillicrap (2015). More recent research has shifted toward developing intelligent exploration techniques Zhang et al. (2021); Ecoffet et al. (2019); Norman & Clune (2023); Mazoure et al. (2023). Methods that decouple exploration and exploitation are particularly relevant to our work Liu et al. (2021); Avner et al. (2012); Norman & Clune (2023); Zhang et al. (2021); Duan et al. (2016); Schäfer et al. (2021); Whitney et al. (2021). In methods, such as First-Explore Norman & Clune (2023) exploration and exploitation are handled by separate policies, with the exploration policy informing subsequent exploitation decisions. Generally, RL focuses on stabilizing the decoupled policies. Evaluations use non-optimal exploitation returns as optimal policies can be hard to obtain in complex environments. Contrary, we measure the optimal exploitation return for the evaluation of LLMs. Agent returns and LLM exploitation do not provide reliable estimates.
@@
Evaluation environments. A variety of environments have been developed to assess the different capabilities of LLMs. Notable examples that require exploration include TextWorld Côté et al. (2019), Minecraft Fan et al. (2022), Crafter Hafner (2021), NetHack Küttler et al. (2020), MiniHack Samvelyan et al. (2021), MiniGrid Chevalier-Boisvert et al. (2024), and BabyAI Chevalier-Boisvert et al. (2018). There is also growing interest in environments for computer and web agents Zhou et al. (2023); Yao et al. (2022a); Xie et al. (2024). Comprehensive benchmarks, including SmartPlay Wu et al. (2023), AgentBench Liu et al. (2023), and BALROG Paglieri et al. (2024), aggregate multiple environments to evaluate agents across a wide range of capabilities. In this context, BALROG reported aimless movements and redundant behaviors. The difference in our work lies in the isolation of exploration. Our experiments are based on mazes and an adapted version of Dark Treasure Room Norman & Clune (2023). Our environments offer limited prior knowledge, provide non-binary, relatively dense rewards, and maintain a high-level action space.
3 Evaluating with Optimal Exploitation
In the following section, we present our evaluation method. We introduce the formal background, the necessity of deriving the optimal exploitation for a fair comparison, and our return decomposition. Additionally, we introduce Maze and Dark Treasure Rooms environments which provide dense rewards and an unsolved challenge for LLM exploration in sequential decision-making.
@@ 3.1 Theoretical Framework
Markov Decision Process. We study a deterministic Markov decision process $(\mathcal{S},\mathcal{A},\mathcal{P},r,\gamma)$ where $\mathcal{S}$ is a finite state space, $\mathcal{A}$ is the finite action space, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ defines the transition dynamics of the environment, $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function and $\gamma$ the discount factor. At each time step $t$, an agent samples an action from policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$ based on the current observation $s_{t}\in\mathcal{S}$ and executes it in the environment. The environment transitions and the agent receives a reward $r_{t}$. In an episodic MDP, the procedure repeats until a terminal state $s_{T}$ is reached or the maximum number of steps exceeded. The sequence of state, action, and reward triples until $s_{T}$ is called a trajectory and describes an episode $\tau_{i}=\left(s_{0},a_{0},r_{0},s_{1},a_{1},r_{1},s_{2},\dots\right)$. In our setting, the agent has access to the current triples and the history of prior trajectories $h_{i}=(\tau_{1},\dots,\tau_{i})$ at each time step $t$ in episode $i$. The goal is to maximize the cumulative return $R_{i}=\sum_{t=1}^{T}r_{t}$.
@@
Exploitation versus exploration. The cumulative return $R^{\text{LLM}}_{i}$ of an agent is influenced by its ability to balance exploitation and exploration. Exploitation leverages past high-rewarding actions to maximize the immediate return based on the agent’s current knowledge. In contrast, exploration emphasizes interacting with the environment to gather new information, which can enhance future decision-making and rewards. An agent fully dedicated to exploitation acts according to policy $\pi_{\text{exploit}}^{*}(h_{i})=\arg\max_{\pi}\sum_{t=0}^{T}r(s_{t},\pi(s_{t}))\text{ with }s_{t},\pi(s_{t})\in h_{i}$ that utilizes past trajectories $h_{i}$ to estimate the optimal action sequence. Conversely, an agent focused on exploration seeks to augment history $h_{i}$ by incorporating new trajectories that maximize the potential future exploitation return. Under the assumption of an optimal exploitation return $R^{\text{exploit}}_{i}$, which ignores redundant or negatively impactful trajectories and whose return monotonically increases over episodes, the objective of the exploration policy is to maximize the exploitation return at the end of exploration: $\pi_{\text{explore}}^{*}=\arg\max_{\pi_{\text{explore}}}\mathbb{E}\left[R^{\text{exploit}}_{\text{final}}\mid h_{i}=h_{i-1}\cup\tau_{\text{explore}}\right]$.
@@
The necessity of optimal exploitation. When an agent does both, exploitation and exploration, progress can be evaluated by temporarily disabling exploration and measuring the cumulative rewards of the agent. However, for an agent fully dedicated to exploration, the return alone does not accurately reflect success (see Figure 2), as successful exploration does not directly result in a higher return. Also measuring exploitation returns with an LLM is similarly hard to exploration and, therefore, does not give reliable estimates (appendix A.7). Nevertheless, in our environments, we can measure the optimal exploitation return $R^{\text{exploit}}_{i}$. Note that this is possible in most evaluation environments by using, for example, dynamic programming methods which guarantee the optimal sequence of actions based on history $h_{i}$. By calculating the optimal exploitation return, we can assess an agent’s exploration capabilities accurately without conflating them with exploitation.
@@
Decomposing the total return gap. Based on the optimal exploitation return, we can decompose the total return gap, i.e., the difference of the maximal achievable return of the environment and the actual achieved return. Let $R^{max}$ denote the maximal achievable reward of the environment; it is independent of the agent and can be predefined for evaluation purposes. As previously, let $R^{LLM}_{i}$ be the agent reward and $R^{exploit}_{i}$ the optimal exploitation reward by acting according to policy $\pi^{*}_{\text{exploit}}$. Contrary to prior evaluations, $\pi^{*}_{\text{exploit}}$ is not determined by an agent. Note that it follows $R^{max}\geq R^{exploit}\geq R^{LLM}$. We let $\Delta^{total}$ denote the total reward gap, which is calculated as the difference of maximal achievable and the actual achieved return by the agent:
@@
The total return gap decomposes into the exploration and exploitation gap.
3.2 Environments
Decision making in room worlds. Our experiments are conducted in interconnected rooms. To mitigate the impact of low-level spatial navigation, we use a higher-level action space that is conceptually similar to TextWorld Côté et al. (2019). Conversely, our domain features only minimal narrative and semantic associations. Each environment has a fixed layout and starting position. Episodes terminate when the agent collects three objects or exceeds a maximum number of door traversals. The latter is calibrated to the distance between the starting position and the furthest room, ensuring that the exploration budget scales with complexity. At the beginning of each episode, the agent is reset to its starting position, requiring it to strategically plan exploration over extended horizons. In our case, the optimal exploitation is calculated by solving an instance of the orienteering optimization problem. More details about the calculation and our environments can be found in appendix A.
Dark Treasure Rooms. The first task follows a grid layout and is an adapted version of Dark Treasure Room Norman & Clune (2023) as we replace the low-level navigation with high-level object interactions. Each map contains balls which are placed randomly on the grid and associated with a fixed reward sampled from a uniform distribution $U(1,10)$. The agent’s goal is to maximize the return by collecting at most three balls within the step budget. As reward locations and their values are randomly generated, the optimal exploration strategy, across multiple layouts, requires full coverage of the state-space. Inadequate exploration risks the exploitation collecting suboptimal rewards due to missed balls. The environments are illustrated in appendix A.1.
Maze. Finally, we compare exploration in randomly generated mazes to scale beyond simple grid layouts. Mazes are similarly common in the RL literature to test in-context learning abilities Laskin et al. (2022); Grigsby et al. (2023); Schmied et al. (2024). Again, we take inspiration from Norman & Clune (2023) and randomly place multiple rewards in the environment. In contrast, to Dark Treasure Rooms, rewards are more sparsely distributed. Further, mazes demand increased long-horizon planning as decisions have more impact on possible future paths. The layouts used for our experiments can be found in appendix A.2 and an exemplary interaction is depicted in Figure 1.